1. Suppose that identical duopoly firms have constant marginal costs of $16 per unit. Firm 1 faces a demand function of q₁ = 70-2p₁ + p₂, where q1 is Firm 1's output, p1 is Firm 1's price, and p2 is Firm 2's price. Similarly, the demand Firm 2 faces is q2 = 70-2p2 + p1. Solve for the Nash-Bertrand equilibrium.


2. Solve for the Nash-Bertrand equilibrium for the firms( described in above question) if both firms have a marginal cost of $0 per unit.


3. Solve for the Nash-Bertrand equilibrium for the firms ( described in above question 1) if Firm 1's marginal cost is $25 per unit and Firm 2's marginal cost is $15 per unit.


1
Expert's answer
2022-01-07T09:19:51-0500

Solution:


a) Solve for the Bertrand equilibrium:


Let c1 represent the marginal cost for Firm 1, and

Let c2 represent the marginal cost for Firm 2.

Here c1=c2=$16.


A Bertrand competitor maximizes profits by choosing the optimal price, taking into account its competitor`s price. For Firm 1, this means:


Π = max{ p1q1(p1,p2)-c1q1}=max{(p1-c1)(70-2p1+p2)}.


Profit maximization implies that "\frac{d}{dp_1}" {(p1-c1)(70-2p1+p2)}=0, or


(70-2p1+p2)-2(p1-c1)=0, or 4p1=70+p2+2c1.


By symmetry, we also have 4p2=70+p1+2c2.


Substituting the second expression into the first gives us "p_1=\frac{1}{4}[70+2c_1+\frac{1}{4}(70+2c_2+p_1)]," or "p_1=\frac{8c_1+350+2c_2}{15}."


By symmetry, "p_2=\frac{8c_1+350+2c_2}{15}."

Substituting the marginal costs above gives:


"p_1=p_2=" $34

"q_1=q_2=" 66


b) Solve for the Bertrand equilibrium if both firms have a marginal cost of $0 per unit:


Substituting the appropriate marginal costs into the expression above gives  


"p_1=p_2=" 23.33

"q_1=q_2=" 46.77


c) Solve for the Nash-Bertrand equilibrium for the firms ( described in above question 1) if Firm 1's marginal cost is $25 per unit and Firm 2's marginal cost is $15 per unit:


Substituting the appropriate marginal costs into the expression above gives :

"p_1=" $38.66

"p_2=" $42.67

"q_1=" $41.33

"q_2=" $65.33